OpenStudy (anonymous):

i need help on this question (cosx)-(cosx)(sin^2x)

geerky42 (geerky42):

HINT: \(\sin^2\theta+\cos^2\theta =1\) so \(\cos^2\theta = 1-\sin^2 \theta\)

hero (hero):

cos(x) is common so factor that out first to get cos(x)(1 - sin^2x)

hero (hero):

From there, you can apply @geerky42's hint

hero (hero):

Actually, what are the instructions for this?

OpenStudy (anonymous):

just to simplify

hero (hero):

Okay, well, everything above is your hint. 1st Factor 2nd Apply Pythagorean Identity 3rd Simplify

OpenStudy (anonymous):

would that come out to -cos^3x?